OpenStudy (anonymous):

A cylindrical vase has a diameter of 4 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inch. The vase is filled with water up to a height of 8 inches. What is the volume of water in the vase?

OpenStudy (anonymous):

26π in3 16π in3 24π in3 32π in3

OpenStudy (valpey):

You need to solve for the total volume of the cylinder, then subtract the total volume of the marbles to get the total volume of the water.

OpenStudy (valpey):

The volume of a cylinder is \(\pi r^2 h\) (r, the radius is half the diameter) The volume of a sphere is \(\frac{4}{3}\pi r^3\)

OpenStudy (anonymous):

oh ok

OpenStudy (valpey):

What is the volume of the cylinder in terms of \(\pi\)?

OpenStudy (anonymous):

well pie is 3.14 right?

OpenStudy (valpey):

It isn't necessary to use the numerical value of pi. You can leave everything in terms of pi. As in \(x = 15\pi\) or \(y = \frac{2}{3}\pi\).

OpenStudy (anonymous):

oh ok well idk what the radius is but they tell us the height

OpenStudy (valpey):

The radius is half the diameter by definition.

OpenStudy (anonymous):

oh ok so the radius would be 2?

OpenStudy (valpey):

Yes, the radius of the cylinder is 2.

OpenStudy (anonymous):

ok so it would be π2^2*8

OpenStudy (valpey):

Yes, what is that equal to?

OpenStudy (anonymous):

315.5072?

OpenStudy (anonymous):

@Valpey

OpenStudy (valpey):

Not quite. Let's still leave it in terms of pi. Let's just resolve \(2^2 * 8\).

OpenStudy (valpey):

I think you did \((\pi*2)^2 * 8\) when you meant to do \(\pi*(2)^2 * 8\). Still, we don't actually need to multiply by pi.

OpenStudy (anonymous):

oh ok 2^2*8 is 32

OpenStudy (anonymous):

and yeah thats what i did